The sum of four consecutive integers is 1344. Find all four integers.
Answer by mananth(16949)

Let the 4 integers be x,x+1,x+2,x+3
..
The sum = 1334 ( I have changed 1344 to 1334)
x+x+1+x+2+x+3=1334
4x+6 = 1334
add -6 to both sides
4x+6-6 =1328-6
4x=1328
divide by 4
4x/4 = 1328/4
x= 332
the integers are 332,333,334,335